A firm plans to invest an amount of money at the beginning of every year in order to acquire a sum of $50,000 at end of 5 years. If the investment rate is 7% p.a., what is the value of the sum


1
Expert's answer
2021-11-16T10:38:56-0500

Future value = Annual payment "\times \frac{(1 + r)^n - 1 }{r } \times (1 + r)"

50000 = Annual payment "\times \frac{(1 + 0.07)^5 -1}{0.07} \times (1+0.07)"

50000 = Annual payment "\times 6.153290741"

Annual payment "= \frac{50000}{6.153290741} = 8125.73"